OpenStudy (anonymous):

3x - 2y + 7x + 4y simplify

OpenStudy (anonymous):

yes plz

OpenStudy (radar):

10x+2y

OpenStudy (turingtest):

10x+2y excuse me 3x-2y+7x+4y combine like terms, i.e. put the x's and y's together and add them with arithmetic 3x+7x+4y-2y=(3+7)x+(4-2)y=10x+2y

OpenStudy (anonymous):

ooooohhh i get it!!